OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). y = 2x + 6 y = −x + 3 y = −x + 6 y = 2x + 3

OpenStudy (primeralph):

NO answer.

OpenStudy (jhannybean):

a line that is perpendicular has a slope that can be found by taking the negative reciproal of the slope from the equation given. m=2 \[\large m_{p}=-\frac{1}{m}\]\[\large m_{p}=-\frac{1}{2}\] your perpendicular slope is -1/2 Now use your points (6,3) into your slope-intercept form and solve for your y-intercept \[\large y=mx+b\]\[\large 3=\left(-\frac12 \right)(6)+b\]\[\large 3 = -3+b\]\[\large b=6\] Now that we've fond our "b" and our slope of our perpendicular line,plug these two points back into slope intercept form. \[\large m_{p}=-\frac12 \ , \ b=6\]\[\large y=mx+b\]

OpenStudy (primeralph):

@Jhannybean Why bother?

OpenStudy (jhannybean):

so the person can learn how to find the perpendicular line and not just look for an answer?

OpenStudy (anonymous):

oh okay I sort of get it now . Thank you very much @Jhannybean , so what 2 answer choican I choose from ?

OpenStudy (jhannybean):

What do you get when you plug in b and m that we've found into the equation?

OpenStudy (anonymous):

It's either first choice or last choice right ?

OpenStudy (anonymous):

So the anwser will be the last choice ? @Jhannybean

OpenStudy (jhannybean):

hm.. m=-1/2 and b =6 y=mx+b y=(-1/2)x+6 Doesn't seem like any of the answer choices. Is this what you got when you plugged in the found values?

OpenStudy (anonymous):

y = 2x + 2 slope for perpendicular line is - 1/2 y = -1/2x + b using points (6,3) 3 = -1/2(6) + b 3 = -3 + b 6 = b y = -1/2x + 6 you got the same thing I got

OpenStudy (anonymous):

So the answer choice will be either a or c right ?

OpenStudy (primeralph):

Wrong input. @justyou_22

OpenStudy (jhannybean):

You seem like you're just guessing, i asked you what you got when you input the values and you havent told me.

OpenStudy (anonymous):

The same thing y'all got y = -1/2 + 6 ?! But that's not an answer choice ??

OpenStudy (primeralph):

@justyou_22 As I said in the first comment, no answer.

OpenStudy (jhannybean):

Exactly @justyou_22

OpenStudy (anonymous):

Oh okay thank you guys (: